# Deep learning framework for action prediction reveals multi-timescale locomotor control

### v4

@@ Abstract
Modeling human movement in real-world tasks is a fundamental goal for motor control, biomechanics, and rehabilitation engineering. However, existing models of essential tasks like locomotion are not applicable across varying terrain, mechanical conditions, and sensory contexts. This is at least in part due to simplifying assumptions like linear and fixed timescales mappings between inputs and future actions, which may not be broadly applicable. Here, we develop a deep learning-based framework for action prediction, outperforming traditional models across multiple contexts (walking and running, treadmill and overground, varying terrains) and input modalities (multiple body states, visual gaze). We find that neural network architectures with flexible input history-dependence, like GRU and Transformer, and with architecture-dependent trial embeddings perform best overall. By quantifying the model’s predictions relative to an autoregressive baseline, we identify context- and modality-dependent timescales. These analyses reveal that there is greater reliance on fast-timescale predictions in complex terrain, gaze predicts future foot placement before body states, and the full-body state predictions precede those by center-of-mass states. This deep learning framework for human action prediction provides quantifiable insights into the control of real-world locomotion and can be extended to other actions, contexts, and populations.
@@ Introduction
Modeling human movement in real-world settings remains a major scientific challenge for motor control, biomechanics, and rehabilitation engineering. Indeed, data-driven models of essential tasks like locomotion have informed predictive simulations [1, 2], helped characterize neuromotor disorders [3, 4], and have been incorporated into wearable robots to anticipate the user’s actions [5, 6]. Data-driven models widely used to characterize locomotor control rely on simplifying assumptions such as linear, fixed-timescale mappings or are limited to laboratory-constrained tasks. Such models have successfully explained how center-of-mass (CoM) kinematics predict foot placement [7, 8, 9, 10] and ground reaction forces [2, 11] during treadmill locomotion, but they have not been shown to generalize to real world settings. During real world movements, people exhibit more variable behaviors [12, 13], and the control strategies may depend on the nonlinear history of multiple task-relevant inputs [14]. Addressing this requires a framework capable of predicting future actions given time-varying inputs from different modalities, such as body state and visual gaze, across contexts. Artificial neural networks have shown significant promise in predicting foot placement actions, as deep learning models can accurately capture the temporal and nonlinear statistics at play [15]. However, previous neural network models of locomotion test a single architecture [16, 17, 18], use a single input modality [16, 17] or are confined to a single laboratory-constrained setting like treadmill walking [18, 19]. Furthermore, these data-driven models do not separate the variance explained by the input modality from the variance explained by the inter-trial variability, hindering model performance and interpretability. Here, we put forth a deep learning framework comparing multiple neural network architectures, with architecture-dependent trial embeddings, outperforming existing models across real world contexts.
Biological motor control is time-varying and context-dependent, requiring adaptive strategies to meet environmental, task-specific, and physiological demands [20]. For example, during locomotion, humans possess a remarkable ability to continuously modulate their actions in diverse “contexts" i.e. environment-specific circumstances such as uneven terrain or obstacle-avoidance [21, 22]. This ability is supported by dynamic sensorimotor control strategies that can be deployed both on a fast timescale such as during each walking step, or on a longer timescale over multiple steps [22, 23]. For instance, walking on uneven terrain requires rapid adjustments to foot placement [24], as well as persistent motor plans involving path or speed selection [25]. In fact, multiple timescales in motor control have been hypothesized to reflect distinct neural computations [26] or distinct functional goals [27], so quantifying multi-timescale control strategies from real world data has theoretical value. Despite this, we lack meaningful approaches to quantify multi-timescale control in real world contexts. Here, we put forth the predictive power of nonlinear action prediction models relative to an autoregressive baseline as a generalizable way to quantify control timescales, discovering a context-dependent tradeoff between fast- and slow-timescale predictions.
Biological sensorimotor control is achieved by dynamically mapping multi-sensory inputs to motor commands, enabling adaptive control in changing contexts [28, 29]. During locomotion, the control of foot placement is inherently multimodal, and prior work has examined the role of individual input modalities such as vision [24] or low-dimensional body states on control [7]. Visual input is crucial for anticipatory locomotor control when navigating obstacles or selecting stable footholds on complex terrains [13, 30], while joint-level or CoM-based feedback informs muscle responses to transient perturbations [31, 32, 2]. Vestibular input is also thought to guide foot placement, particularly in a phase-dependent manner [33]. Despite the existence of multiple input modalities, existing foot placement control models assume a linear fixed control timescale with low-dimensional body states as an input [7, 34], thereby neglecting the temporal dynamics of how and when each modality contributes to control, and how different modalities compare to one another. To address these gaps, we propose a data-driven framework that systematically evaluates the contributions of multiple input modalities — including CoM-relevant kinematics, full-body kinematics, and gaze fixations — to predicting future foot placement. By comparing the predictive power of each modality to an autoregressive baseline model, we quantify the timescale and magnitude of modality-dependent foot placement. Our framework reveals at which phase of the gait cycle each input modality become critical for informing future foot placement.
In this study, we develop deep learning-based models to predict future actions from multimodal input time series and test them in diverse locomotor tasks. By comparing different model architectures, we find that architectures with flexible input history-dependence generalize better across real-world contexts and input modalities. We characterize the control timescale of a given nonlinear action prediction model using its predictive power relative to an autoregressive baseline model, to help reveal how these timescales are shaped by the context and the input modality. Our work has implications for characterizing, modeling, and predicting real world movements and the multi-timescale control strategies hidden in them. This deep learning framework can inform the development of human-aware wearable robot control and more human-like physics-based simulations of motor behavior in the real world [35, 36].
@@ Results
In this section, we present findings using our deep learning framework for action prediction to understand the influence of the environmental context and input modality on future foot placement. First, we demonstrate that neural network models, particularly GRUs, significantly outperform linear models in more challenging contexts. During overground walking, we find that subjects exhibit a tradeoff between fast and slow timescale predictions, with the magnitude of the tradeoff varying across different terrains. By comparing the predictive power of input modalities to an autoregressive baseline model, we reveal that control timescales are both modality- and context-dependent. Finally, we discover a gait phase mid-swing when the swing foot dynamics becomes useful for foot placement prediction. We encourage readers to first read through the Methods section to understand the terminology used below. An overview of the key methodological components in the deep learning framework can be found in Figure 1.
@@
Figure 1: Overview of the deep learning framework. A We include diverse tasks and contexts, including overground walking with various terrain conditions, treadmill walking, and treadmill running on even and uneven terrain. B Inputs contain hypothesized modalities such as CoM-relevant kinematics, swing-foot kinematics, full-body kinematics, and gaze fixations (defined by the intersection of the gaze vector with the ground plane). The output is foot placement. C Foot placement output is defined by the lateral and fore-aft foot placement relative to the opposite foot at its previous heel strike. D The input is a time series spanning from three strides prior to touchdown to a specific gait phase within the previous stride. The output corresponds to foot placement at heel strike. We denote the ending gait phase of the time series relative to one stride prior to touchdown as $\phi$, discretized and sampled from $n=21$ equally spaced phases within $0\leq\phi\leq 1$. E The network architecture of the nonlinear models. The input consists of a time series, an L/R flag, and a trial ID. The time series and L/R flag are concatenated, while the trial ID is embedded to provide additional trial-specific information. The incorporation of this embedding is architecture-dependent, i.e. it influences the model prediction differently depending on the neural network architecture, as detailed in the Methods section.
@@
Models with flexible nonlinear input dependence outperform linear models
@@
We evaluated the performance of nine models — LSTM, GRU, FCNN, TCN, Transformer, Linear instance with (LI2) and without (LI) L2 regularization, Linear history with (LH2) and without (LH) L2 regularization — on their ability to predict foot placement across various tasks and input modalities. The tasks included treadmill walking, treadmill running, and overground walking on various terrains, while the input modalities consisted of CoM-relevant kinematics, full-body kinematics, swing-foot kinematics and gaze fixations. The model score was evaluated based on the Root Mean Squared Error (RMSE) relative to the optimal model at each gait phase, and then normalized by the highest-scoring model. The Methods section details the evaluation metric.
Task Terrain M LSTM GRU FCNN TCN Trans LI LH LI2 LH2
TW N/A C 1.00 1.00 0.98 0.71 0.98 0.97 0.26 0.97 0.99
C 0.98 1.00 0.96 0.92 0.98 1.00 0.34 1.00 0.97
Even F 1.00 1.00 0.89 0.82 0.94 0.38 0.62 1.00 0.78
C 0.98 1.00 0.96 0.91 0.97 0.97 0.31 0.97 0.97
TR Uneven F 1.00 1.00 0.93 0.91 0.96 0.35 0.63 0.99 0.76
C 0.96 1.00 0.95 0.94 0.99 0.91 0.70 0.95 0.88
F 0.97 0.99 0.94 0.95 1.00 0.52 0.76 0.87 0.77
Flat V 0.80 0.87 0.70 0.91 1.00 0.69 0.30 0.75 0.30
C 0.96 0.99 0.95 0.98 1.00 0.86 0.66 0.93 0.78
F 0.97 0.98 0.94 0.99 1.00 0.52 0.76 0.82 0.77
Medium V 0.95 0.98 0.88 1.00 0.97 0.68 0.32 0.75 0.33
C 0.99 1.00 0.97 0.99 1.00 0.31 0.68 0.87 0.73
F 0.99 0.99 0.94 0.98 1.00 0.58 0.74 0.74 0.74
OW Rough V 1.00 1.00 0.92 0.96 0.97 0.34 0.61 0.61 0.61
Table 1: Normalized Model Scores across Tasks and Input Modalities. Models were evaluated on treadmill walking (TW), treadmill running (TR), and overground walking (OW) tasks with varying terrain. The input modalities (M) were: CoM-relevant kinematics (C), full-body kinematics (F), and visual gaze (V). The normalized model scores use the following color coding: dark green (0.980.980.980.981.001.001.001.00), light green (0.950.950.950.950.980.980.980.98), orange (0.900.900.900.900.950.950.950.95), and red ($<0.9$). Normalized model scores, defined in the Methods section, are rounded to two decimal places, potentially resulting in two models having a score of $1.00$ (e.g., the other model’s score could be $0.997$).
@@
Our analysis reveals that during treadmill walking and running, the locomotor control strategies can be effectively captured by linear models, particularly model LI2. Among the tested architectures, GRU demonstrates the best overall performance, followed closely by Transformers and LSTM (Table 1). Since GRU demonstrates the best overall performance, it is used for the analyses in the subsequent sections. An exception arises when using gaze fixations as the input modality during flat terrain overground walking, where Transformers significantly outperform all other models. However, this is also the only modality and context where predictive power never surpasses the baseline model, suggesting that gaze does not play a significant role in foot placement prediction. To examine the temporal trends of each model more closely, we provide the RMSE as a function of gait phases for each model under different conditions in Figures 6 and 7. We find that FCNN performs better during treadmill locomotion, but poorly during overground walking, whereas TCN performs poorly during treadmill locomotion, even worse than some linear models, but better during overground walking. While linear models are sufficient to capture the relationship between input modality and foot placement during treadmill movements, nonlinear models are necessary to account for the history-dependence and inherent nonlinearities of the input modality in non-stationary contexts (Table 1).
Fast- versus slow-timescale predictions are context-dependent
Biological motor control exhibits remarkable adaptability, seamlessly combining fast- and slow-timescale control strategies, depending on task complexity and other environmental demands. Here, we quantify multi-timescale tradeoffs as a function of varying terrains. We define the tradeoff between fast- and slow-timescale predictions as the relationship between CoM-relevant prediction intercept and its peak predictive power relative to the autoregressive baseline. The intercept is defined by the $R^{2}$ value at gait phase 0 (Figure 2A), quantifying the degree of long-timescale prediction. We obtain the relative predictive power ($\Delta R^{2}$) by subtracting the baseline predictive power from the $R^{2}$, and define the peak relative predictive power as maximum $\Delta R^{2}$. While trials exhibit varying intercepts, the intercepts of CoM-relevant kinematics and the baseline are highly correlated (Figure 2B; $r=0.97$). By subtracting baseline predictive power from predictive power, this relative measure reduces inter-trial variability, enabling a more meaningful comparison across trials.
During overground walking on flat terrain, the negative correlation reveals a tradeoff between long- and short- timescale prediction, indicating that trials that rely more heavily on long-timescale prediction shows a reduced dependency on short-timescale prediction (Figure 2C, $r=-0.69$). As terrain becomes rougher, subjects exhibit less long-timescale prediction, leading to a decrease in the intercept of CoM-relevant kinematics, thereby reducing the magnitude of the correlation with the peak relative predictive power (Figure 2C; $r=-0.5$ on medium terrain, $R=0.02$ on rough terrain). Further analysis of the slope between the peak of the relative predictive power and the intercept reveals additional insights; the negative sign of the slope confirms a tradeoff between long- and short-timescale predictions, while the slope magnitude represents the rate at which long-timescale prediction decreases as short-timescale prediction increases. We observe that the tradeoff magnitude changes across terrain variations, reflecting a change in long- and short-timescale prediction of foot placement (Figure 2D).
@@
Figure 2: Relative predictive power reveals fast and slow timescales across contexts. A Explained variance ($R^{2}$) as a function of gait phase for modality-based (orange) and autoregressive baseline (blue) predictions. The intercept is defined as the modality-based $R^{2}$ value at gait phase 0 (i.e., the input time series spans from three gait cycles prior to touchdown and ends one gait cycle before touchdown). The red shaded area represents the relative predictive power ($\Delta R^{2}$), calculated as the difference between modality-based and baseline predictions. The peak is the maximum $\Delta R^{2}$. B Baseline intercept versus CoM-relevant kinematics intercepts for treadmill walking. C Relationship between the CoM-relevant kinematics intercepts and peak $\Delta R^{2}$ during overground walking across terrains with linear regression fits. The plots display the regression line and 95% confidence interval. D Coefficients between CoM-relevant kinematics $\Delta R^{2}$ peak and intercept of CoM-relevant kinematics. The box-plot shows the median (bar in box), $25-75\%$ percentile (box), and non-outlier range (whiskers) of the bootstrap statistics. The magnitude of the slope quantifies the strength of the tradeoff between long- and short-timescale prediction.
Relative predictive power reveals different control timescales across modalities
Human locomotion involves dynamic integration of multiple sensory inputs to guide future foot placement. Our framework seeks to uncover how humans utilize various input modalities, including CoM-relevant kinematics, full-body kinematics, and gaze fixations, across varying contexts like treadmill walking [7], treadmill running [37], and overground walking on terrains of different roughness [30]. To understand the predictive ability of the input modality, we train a model that takes the time series of the input modality and predicts the future foot placement. As the input time window increases, the model demonstrates improved predictive power, as shown by a closer alignment between predictions and ground truth (Figure 3A). The results demonstrate how different input modalities influence foot placement prediction across contexts. By analyzing the relative predictive power ($\Delta R^{2}$) of modality-based models compared to the autoregressive baseline, we obtain the timing and magnitude of each modality’s contribution to foot placement control (Figure 3B).
During overground walking on flat terrain, gaze does not play a significant role in foot placement prediction, as its performance does not surpass the autoregressive baseline (Figure 3C). This suggests that foot placement control on flat terrain is not guided by gaze. Across contexts, full-body kinematics begin to outperform the baseline earlier than CoM-relevant kinematics; we provide some possible explanations in the discussion section. Despite the inclusion of significantly more body markers (29 vs. 1), full-body kinematics only explained 6-14% of additional foot placement variance, which highlights the significance of CoM-relevant kinematics in locomotor control (Figure 3C). As terrain complexity increases, foot placements are better predicted by gaze, with predictions earlier than both CoM-relevant and full-body kinematics for lateral (Figure 3C) but not fore-aft (Figure 8C) adjustments. This indicates that in more challenging environments, humans rely on visual gaze to inform lateral foot placement in advance of the body states. Across contexts and modalities, lateral foot placement control occurs earlier than fore-aft control, indicating a prioritization of lateral stability during locomotion (Figure 8A), which aligns with previous findings [7, 38, 39]. During treadmill running, the relative predictive power for lateral foot placement is not significantly different between even and uneven terrain. However, the fore-aft foot placement predictions is earlier during on uneven than on even terrain (Figure 8B). This discrepancy is likely a result of solely fore-aft terrain perturbations introduced in the uneven terrain running study [37].
Figure 3: Relative predictive power reveals modality-dependent control timescales. A Relative predictive power ($\Delta\,R^{2}$, median and interquartile range) of CoM-relevant kinematics, full-body kinematics, and gaze fixations for overground walking on different terrain roughness. The highlighted interval begins when $\Delta\,R^{2}$ exceeds 5% (Wilcoxon, one-sided), and ends at the peak $\Delta\,R^{2}$. The shaded area represents the swing phase. B Relative predictive power of CoM-relevant kinematics for treadmill walking. C Relative predictive power of CoM-relevant kinematics and full-body kinematics for treadmill running on even and uneven terrains.
The swing foot dynamics predict future foot placement mid-swing
Swing foot dynamics are essential for understanding how humans adapt their locomotion strategies to varying terrains and tasks. In prior work, it is implicitly assumed that swing foot prediction occurs simultaneously with the timing of swing initiation [11, 40]. To analyze these dynamics, we detect the timing of swing initiation by identifying when the swing foot’s velocity in the direction of locomotion exceeds 5% of its peak velocity (Figure 5B). During overground walking, swing initiation occurs significantly later than during treadmill walking (Wilcoxon signed-rank test, one-sided, $p<0.05$; Figure 4A). As the terrain becomes rougher, the timing of swing initiation is further delayed, perhaps reflecting the need to gather additional information for precise foot placement on a stable foothold. In contrast, running exhibits different swing foot prediction dynamics [41], characterized by the absence of a double stance phase and the presence of a flight phase, where neither foot is in contact with the ground. These features allow for much earlier swing initiation. Similarly, during treadmill running on uneven terrain, swing initiation occurs later, consistent with the increased demands of navigating complex environments.

### v5

# Human locomotor control timescales depend on the environmental context and sensory input modality

@@ Abstract
Everyday locomotion is a complex sensorimotor process that can unfold over multiple timescales, from long-term path planning to rapid, reactive adjustments. However, we lack an understanding of how factors such as environmental demands, or the available sensory information simultaneously influence these control timescales. To address this, we present a unified data-driven framework to quantify the control timescales by identifying how early we can predict future actions from past inputs. We apply this framework across tasks including walking and running, environmental contexts including treadmill, overground, and varied terrains, and sensory input modalities including gaze fixations and body states. We find that deep neural network architectures that effectively handle long-range dependencies, specifically Gated Recurrent Units and Transformers, outperform other architectures and widely used linear models when predicting future actions. Our framework reveals the factors that influence locomotor foot placement control timescales. Across environmental contexts, we discover that humans rely more on fast timescale control in more complex terrain. Across input modalities, we find a hierarchy of control timescales where gaze predicts foot placement before full-body states, which predict before center-of-mass states. Our model also identifies mid-swing as a critical phase when the swing foot’s state predicts its future placement, with this timescale adapting across environments. Overall, this work offers data-driven insights into locomotor control in everyday settings, offering models that can be integrated with rehabilitation technologies and movement simulations to improve their applicability in everyday settings.
@@ Introduction
Understanding the mechanisms humans use to control everyday movements is a fundamental scientific challenge for sensorimotor control, neuromechanics, and rehabilitation engineering. For instance, a critical aspect of locomotor stability is the control of foot placement relative to the body [1, 2]; the timescale on which this control unfolds can vary from faster reactive control [3, 4] to slower proactive selection of footholds [5, 6]. Moreover, the existence of multiple control timescales likely maps to distinct sensorimotor and neural underpinnings [7, 8], and may serve different environment-dependent functional goals [9, 10]. To achieve adaptive locomotion, humans must therefore modulate control across multiple timescales informed by the environmental context and the available sensory information. Despite advances in collecting high-throughput movement data in real-world settings, we lack an understanding of how multiple control timescales are chosen and how they depend on the environmental and sensory context. This study introduces a data-driven framework to characterize multi-timescale control strategies in locomotion across diverse environmental settings and sensory input modalities.
Figure 1: A Humans integrate sensory inputs to plan for future foot placements during locomotion. While sensory inputs such as visual, proprioceptive and vestibular information are not readily measurable in humans, we use gaze fixation, full-body and center-of-mass kinematics to indirectly infer their predictive influence on foot placement. B We analyze walking data across varied terrains to explore how humans adapt foot placement control strategies in response to increasing environmental complexity. The heatmaps illustrate lateral and fore-aft right foot placement relative to the left foot across terrains, with gradients derived from kernel density estimation. C We seek to quantifies the control timescale under diverse environmental contexts, and across various input modalities, including gaze fixation, full-body, center-of-mass, and swing foot kinematics.
Using data-driven models to understand multi-timescale control strategies in complex human behaviors necessitates some form of “system identification” i.e., mapping future actions to a history of time-varying inputs. While early data-driven models for system identification of locomotor control assume linear and fixed timescale mappings between states and actions [11, 12, 13], it is unclear whether these assumptions generalize to more complex settings. Data-driven models that leverage deep learning architectures offer a promising approach for capturing such context-dependent and nonlinear dynamics [14]. However, prior attempts to develop such nonlinear data-driven models of locomotor control have not been used to understand the underlying control strategies and have been limited to a single network architecture [15, 16], a single input modality [17, 15], or to laboratory-constrained datasets [16, 18]. To overcome these limitations, we put forward a generalizable data-driven framework to understand locomotor control in real-world settings, testing its predictive performance across different network architectures, different input modalities (e.g. gaze, full-body kinematics), and multiple environmental contexts (e.g. varied terrains). Using these models, we provide insight into how the environmental context and sensory input modality influence multi-timescale control.
Biological motor control integrates information from multiple sensory input modalities to plan future actions [19, 20]. During locomotion, foot placement is influenced by vision [21, 5], center-of-mass (CoM) states [11, 12], and postural information [22, 23]. While direct measurement of the underlying sensory inputs (e.g. visual, proprioceptive, and vestibular signals) is typically impossible in humans, we can investigate their influence by studying measurable proxies like gaze fixations, full-body states, and CoM states (Figure 1A). Existing data-driven models of locomotion often oversimplify the influence of multiple sensory input modalities by focusing on a single input and assuming a fixed control timescale [11, 13]. In this study, we systematically evaluate how the timescale of foot placement control is influenced by various input modalities by characterizing when information from a given input becomes available for prediction (Figure 1C). By characterizing how the control timescales depend on the sensory input modality, we reveal critical phases within the movement when each input becomes useful for control.
Biological motor control is continuously tuned to environment-specific demands, employing adaptive strategies [24]. To maintain stability, individuals adjust their actions in response to changing environmental features like varying terrains or obstacles [25, 26]. As the environmental complexity increases, humans may adjust their sensorimotor control strategies (Figure 1B); perhaps by relying more on visual gaze [27, 25] than on other sensory input modalities, thereby explicitly prioritizing safer footholds over more automatic foot placement choices. These adjustments can occur over multiple timescales, ranging from within-step corrections to longer-term strategic planning over several steps [26, 28, 6]. The possibility of multiple control timescales could also point to distinct underlying neural computations [7] or distinct functional goals [9, 10]. It is therefore a necessary first step to quantify these control timescales to advance our scientific understanding of motor control in real-world settings. Here, we use the relative predictive power of data-driven models trained on distinct input modalities to identify these control timescales, thereby revealing how they depend on the environmental context. Specifically, we discover an environmental context-dependent shift from slow to fast control timescales in human locomotion.
In this study, we (i) develop data-driven models, leveraging deep learning architectures, to predict future foot placements from past input states, and (ii) utilize these models to understand how the timescale of locomotor control depends on the environmental context and the sensory input modality. We found that neural network architectures suited to capturing long-range dependencies generalize best across different contexts. Our data-driven models reveal that the locomotor control timescale depends on both the environmental context and the sensory input modality, highlighting when different inputs become useful for control and how the control timescale shifts with environmental demands. Collectively, this work offers data-driven insights into the influence of environmental context and sensory information on locomotor control, and these data-driven models can be extended to characterize sensorimotor disorders, design human-aware wearable assistive technologies, and develop more human-like simulations of complex everyday human movements.
@@ Results
In this section, we present findings developing a data-driven framework to predict future actions from past input states and using these models to understand how the environment and the input modality influence locomotor foot placement control timescales. First, we find that deep learning architectures, particularly GRUs and Transformers, significantly outperform other architectures and widely used linear models across different environmental settings. Using these models, we find that the degree to which humans rely on faster versus slower foot placement predictions during overground walking depends on the terrain complexity. Next, we discover that the swing foot input state becomes predictive of future foot placement in the middle of the swing phase, and how this timing changes across terrains. Finally, by comparing the predictive power of distinct input modalities to a baseline, we characterize how the timescale of prediction depends on the sensory input modality. An overview of the key methodological components of the data-driven framework can be found in Figure 2.
@@
Figure 2: Overview of the data-driven framework. A We include diverse tasks and contexts, including overground walking with various terrain conditions, treadmill walking, and treadmill running on even and uneven terrain. B Inputs contain hypothesized modalities such as CoM-relevant kinematics, swing foot kinematics, full-body kinematics, and gaze fixations (defined by the intersection of the gaze vector with the ground plane). The output is foot placement. C We define foot placement output as the lateral and fore-aft foot placement relative to the opposite foot at its previous heel-strike. D We use a time series spanning from three strides prior to touchdown to a specific gait phase within the previous stride. The output corresponds to foot placement at heel-strike. We denote the ending gait phase of the time series relative to one stride prior to touchdown as $\phi$, discretized and sampled from $n=21$ equally spaced phases within $0\leq\phi\leq 1$. E The network architecture of the nonlinear models. The input consists of a time series, an L/R flag, and a trial ID. The time series and L/R flag are concatenated, while the trial ID is embedded to help the model learn trial-specific information. We incorporate this embedding in an architecture-dependent manner, i.e. it influences the model prediction differently depending on the neural network architecture, as detailed in the Methods section.
@@
Models with flexible nonlinear input dependence best predict future foot placement across settings
@@
We aimed to test which data-driven model architectures best predict future foot placement from a history of input states, comparing many deep neural network architectures and traditionally used linear models [11, 29] across input modalities and environmental contexts. Specifically, we evaluated the performance of nine models — GRU, Transformer, LSTM, FCNN, TCN, Linear instance with (LI2) and without (LI) L2 regularization, Linear history with (LH2) and without (LH) L2 regularization (see Methods for details) — based on their ability to predict foot placement across settings including treadmill walking, treadmill running, and overground walking on various terrains, and input modalities including CoM-relevant kinematics, full-body kinematics, swing foot kinematics, and gaze fixations. We consider the predictive power of the model using swing foot kinematics as the baseline model, which indicates how much information is contained in the foot itself over the gait cycle for predicting future foot placement. We evaluated each model’s performance based on the Root Mean Squared Error (RMSE) of its predictions relative to that of the optimal model at each gait phase, and then normalized by the highest-scoring model. The Methods section details the baseline model and the evaluation metric.
@@
We find that during treadmill walking and running, the body state to foot placement mappings are effectively captured by linear models, particularly model LI2 which includes regularization (within 3% of performance; Table 1). However, the linear models are severely outperformed by nonlinear models in the overground locomotion tasks (5-13% worse for CoM as input modality, 13-39% worse for other input modalities; Table 1). Nonlinear models are therefore essential for capturing the history-dependence and inherent nonlinearities that map input states to foot placements during non-stationary activities, such as walking overground. Among the nonlinear architectures, GRU demonstrates the best overall performance across tasks, followed closely by Transformer and LSTM architectures. An exception arises when using gaze fixations as the input modality during flat terrain overground walking, where Transformers significantly outperform all other models. However, this is also the only input modality and context where the predictive power of the model with the gaze as input to predict future foot placement never surpasses the baseline model which uses the foot state to predict future foot placement, suggesting that gaze does not play a significant role in foot placement prediction during overground walking on flat terrain. To examine the temporal trends of each model more closely, we provide the RMSE as a function of gait phases for each model under different conditions in Figures 7 and 8 in the Supplementary Information. We find that FCNN performs better during treadmill locomotion, but poorly during overground walking, whereas TCN performs poorly during treadmill locomotion, even worse than the optimal linear model, but better during overground walking. Since GRU demonstrates the best overall performance, we use it for the analyses in the subsequent sections of the Results.
Task Terrain M GRU Transformer LSTM FCNN TCN LI LH LI2 LH2
TW N/A C 1.00 0.98 1.00 0.98 0.71 0.97 0.26 0.97 0.99
C 1.00 0.98 0.98 0.96 0.92 1.00 0.34 1.00 0.97
Even F 1.00 0.94 1.00 0.89 0.82 0.38 0.62 1.00 0.78
C 1.00 0.97 0.98 0.96 0.91 0.97 0.31 0.97 0.97
TR Uneven F 1.00 0.96 1.00 0.93 0.91 0.35 0.63 0.99 0.76
C 1.00 0.99 0.96 0.95 0.94 0.91 0.70 0.95 0.88
F 0.99 1.00 0.97 0.94 0.95 0.52 0.76 0.87 0.77
Flat V 0.87 1.00 0.80 0.70 0.91 0.69 0.30 0.75 0.30
C 0.99 1.00 0.96 0.95 0.98 0.86 0.66 0.93 0.78
F 0.98 1.00 0.97 0.94 0.99 0.52 0.76 0.82 0.77
Medium V 0.98 0.97 0.95 0.88 1.00 0.68 0.32 0.75 0.33
C 1.00 1.00 0.99 0.97 0.99 0.31 0.68 0.87 0.73
F 0.99 1.00 0.99 0.94 0.98 0.58 0.74 0.74 0.74
OW Rough V 1.00 0.97 1.00 0.92 0.96 0.34 0.61 0.61 0.61
Table 1: Normalized model scores across tasks and input modalities. Models were evaluated on treadmill walking (TW), treadmill running (TR), and overground walking (OW) tasks with varying terrain. The input modalities (M) were: CoM-relevant kinematics (C), full-body kinematics (F), and visual gaze (V). The normalized model scores use the following color coding: dark green (0.980.981.001.00), light green (0.950.950.980.98), orange (0.900.900.950.95), and red (<0.9<0.9). Normalized model scores, defined in the Methods section, are rounded to two decimal places, potentially resulting in two models having a score of $1.00$ (e.g., the other model’s score could be $0.997$).
Faster versus slower prediction timescales depend on the environmental context
Biological motor control exhibits remarkable adaptability, seamlessly combining fast and slow control strategies to meet environmental demands. Here, we quantify these timescales by measuring how well an individual’s body states predict their future foot placement across different environments (Figure 1), using the data-driven framework from the previous section. For clarity, we first outline how the faster and slower timescales are computed (also shown in Figure 3B):
  • We define the slower prediction timescale through the intercept of a model trained on the CoM kinematics. This intercept is the $R^{2}$ of the model at gait phase 0, i.e. using an input window spanning from six to two steps before foot placement.
  • We define the faster prediction timescale through the peak of the relative predictive power of the CoM-based model (Figure 3B). This relative power ($\Delta R^{2}$) is calculated by subtracting the predictive power of a baseline model (which predicts foot placement from past foot states) from the CoM-based model $R^{2}$. The peak is the maximum of this $\Delta R^{2}$.
While different trials show varying intercepts, the intercepts of the CoM-based kinematics and the baseline model are highly correlated ($r=0.97$). Thus, subtracting the baseline predictive power from the CoM predictive power effectively reduces inter-trial variability, providing more interpretable comparisons across environmental contexts. We found a trade-off between these fast and slow control timescales during overground walking, revealed by a negative correlation that depends on the environment. This negative correlation indicates a greater reliance on either slow- or fast-timescale prediction depending on the terrain (Figure 3A, inset). As the terrain becomes rougher, individuals rely less on slow-timescale prediction, indicated by a decrease in the intercept of the CoM-based kinematics and a decrease in its correlation with the $\Delta R^{2}$ peak, from r=0.69r=-0.69 on flat terrain to r=0.5r=-0.5 on medium terrain, and $r=0.02$ on rough terrain (Figure 3A). The change in the slope of the relationship between peak relative predictive power and the intercept across terrains suggests that the balance between slow- and fast-timescale foot placement predictions shifts with the environmental context (Figure 3C).
@@
Figure 3: Relative predictive power reveals fast and slow timescales across contexts. A Relationship between the CoM-relevant kinematics intercepts and peak $\Delta R^{2}$ (both defined in B) during overground walking across terrains with linear regression fits. The plots display the regression line and 95% confidence interval. B Explained variance ($R^{2}$) as a function of gait phase for modality-based (orange) and autoregressive baseline (blue) predictions. The intercept is defined as the modality-based $R^{2}$ value at gait phase 0 (i.e., the input time series spans from three gait cycles prior to touchdown and ends one gait cycle before touchdown). The red shaded area represents the relative predictive power ($\Delta R^{2}$), calculated as the difference between modality-based and baseline predictions. The peak is the maximum $\Delta R^{2}$. C Linear coefficients between CoM-relevant kinematics $\Delta R^{2}$ peak and intercept of CoM-relevant kinematics. The box-plot shows the median (bar in box), $25-75\%$ percentile (box), and non-outlier range (whiskers) of the bootstrap statistics. The magnitude of the slope quantifies the strength of the tradeoff between fast- and slow-timescale prediction.
Swing foot kinematics predict future foot placement mid-swing